# Bi-directional Model Cascading with Proxy Confidence

### v1

@@ Abstract.
Machine Learning, NLP, Model Cascades
ccs: Computing methodologies Machine learningccs: Computing methodologies Neural networks
1. Introduction
@@
In the last decade large, transformer-based, pretrained language models (LLMs) have surged in both popularity and capability, radically improving the state of the art in a wide range of tasks, from traditional Natural Language Processing (NLP) tasks to more general problems that were previously impossible, such as complex logical reasoning and the application of expert knowledge. However, these abilities have often required increasingly large model sizes where inference is expensive, raising concerns ranging from user budgets (Chen et al., 2023b) to environmental impacts (Singh et al., 2024; Ren et al., 2024), including among the NLP community (Arase et al., 2021). In response to this, a diverse range of techniques have been developed to accelerate or improve the efficiency of LLMs.
@@
Some approaches aim to reduce the size of individual models like quantization (Yao et al., 2022) and model pruning (Sun et al., 2024). Others have drawn from the field of adaptive computation, developing systems of models that collaborate to improve some aspect of performance. Conceptually relying on the fact that not all tasks are equally difficult, these approaches make the inference process dependent on the individual sample, leveraging smaller models or subsystems to handle easy instances. This concept has been used to provide large improvements in model throughput using draft-and-verify paradigms (Chen et al., 2023a; Leviathan et al., 2023) as well as to optimize the accuracy-efficiency trade-off (Nie et al., 2024). Early-exit model cascading is a technique that has been used in other machine learning domains for some time (Viola and Jones, 2001; Park et al., 2015) that takes advantage of this principle by engaging a sequence of models of increasing size and capability. Inference is performed on each sample starting with the smallest model and exiting the cascade if the output of any model appears sufficient by some measure of confidence. This approach has recently been applied to NLP and LLMs and has delivered significant efficiency improvements (Yue et al., 2024). However, there are some notable limitations that we address in this paper.
@@
In cascading models, there is typically a sequence of models of increasing complexity, $M_{1}\ldots M_{K}$, and Jitkrittum et al. (2023) observe that standardly, the decision to defer from a smaller model $M_{i}$ to larger model $M_{i+1}$ is based only on some measure of confidence from the smaller model $M_{i}$. This can result in various failure modes, for example in confidence-based deferral erroneously forwarding samples where $M_{i+1}$ performs worse than $M_{i}$. Jitkrittum et al. (2023) carry out a theoretical analysis for the two-model case that defines oracles that include probabilities from the larger model, formally characterising those cases where deferrals based on small-model probabilities alone are insufficient or non-optimal. For practical use, they define post-hoc estimators of their oracle rules that include theoretical confidence of both small and large models. Importantly, they note that since querying the larger model defeats the entire purpose of cascades, they use only the outputs of the small model in their estimators. Their experiments on computer vision datasets, however, do not show much improvement over the standard use of small model confidence.
@@
The first core idea of the present paper is to use large model confidences obtained without querying the model, but instead by building a much smaller proxy model whose only role is to predict large model confidences. For this, we draw on recent work on calibrating LLMs by Ulmer et al. (2024) that will allow us to use only model inputs. We then have ‘backward’, post-invocation confidence estimation for the smaller model, and a ‘forward’, pre-invocation confidence estimation for larger models using a small auxiliary model.
@@
There is then the question of how to relate these two kinds of confidence estimation. We cannot expect that they will be calibrated with respect to each other: in the first place, one is directly taken from the inference model and the other from a proxy; in addition, it is known that larger models produce better calibrated results than smaller models (Zhu et al., 2023). Our second core idea is then to use a meta-model for making deferral decisions in a cascade, extending their use from single-model calibration (Chen et al., 2019; Shen et al., 2023). This then requires a richer representation of features for model confidence: where existing cascading work largely uses final-layer maximum softmax probability output or entropy, we use intermediate layer probing results, which have been shown to be effective in improving calibration in various contexts (Papernot and McDaniel, 2018; Mielke et al., 2022; Chuang et al., 2024).
@@
  • We present a new bi-directional approach to cascading based on a meta-model that incorporates (i) richer representations of multiple model confidences that use internal model states, and (ii) novel proxy confidence representations that avoid querying larger models.
  • We find that the bi-directional approach outperforms methods that use only smaller model confidence. We also find that, by itself, using internal state representations for smaller model confidence improves over the standard use of final layer maximum probability.
@@ 2. Related Work
The accuracy-efficiency trade-off is a pervasive concern in ML, particularly in the field of LLMs, and there are many methods for reducing the computational requirements of models while minimally compromising on output quality (Kaplan et al., 2020). The idea of making an early-exit from inference is one that has existed for many years (Viola and Jones, 2001), but has seen renewed interest with the increasing size and commercial viability of modern models (Rahmath P et al., 2024). Broadly this involves managing the interaction between each sample and the inference system to minimize wasteful computation. This can been applied at the model-level, directing the flow of samples through a set of discrete models, or at the layer-level, controlling the number of dependent layers to be used within a single neural network model (Miao et al., 2024).
2.1. Model Cascades
@@
In early-exit model cascades inference is first performed on a sample using the smallest available model, then some estimate of the confidence of the result is used to determine whether the sample should be deferred to the next largest model or the output should be accepted as is. This process occurs iteratively until the output is accepted or the largest model in the cascade is invoked (Varshney and Baral, 2022). The decision to accept or defer is typically a comparison between the confidence and some static or dynamic threshold, which can be adjusted to satisfy the desired trade-off (Wang et al., 2017; Lebovitz et al., 2023).
@@
Approaches in existing literature vary in number and type of models in the cascade. Lebovitz et al. (2023) find that two model cascades perform well, with three model cascades performing marginally better at the cost of increased system complexity, while others (Chen et al., 2023b) include many models to take advantage of disparate commercial API prices. Nie et al. (2024) cascade 3 heterogeneous models (Logistic Regression, BERT and GPT/LLaMA) with online optimization, treating the largest model outputs as ground-truth with strong results; however, this constrains the application to tasks where traditional models (LR) are viable and assumes the success of largest available model. Choice of confidence metric $A$ also varies, with maximum probability and entropy both being frequently used and effective with some conflicting results regarding which is better (Lebovitz et al., 2023; Wang et al., 2017). Our work broadly fits into this part of the space, following the observations of Jitkrittum et al. (2023) about the potential (but not necessarily realised) value of using confidence from multiple models in a cascade.
@@
Other recent work has focussed on LLM-specific features. To note three: Gupta et al. (2024) use language model cascades to accomplish generative tasks; Yue et al. (2024) use Chain-of-Thought prompting to improve small model confidence; (Wang et al., 2024) alter the LLM training paradigm to produce models best suited to cascading.
2.2. LLM Calibration
@@
A model’s calibration is its ability to express some level of confidence associated with an output that accurately reflects the likelihood that it is correct (Guo et al., 2017). LLM calibration generally has great value to researchers and industry applications, and is the subject of significant research efforts (Geng et al., 2024). The acceptance of LLMs in real-world application depends on their trustworthiness and reliability (Bhatt et al., 2021), while in cascading context, calibration has direct and measurable benefits on performance. The success of any model cascade depends on the quality of its deferral process, and therefore the calibration of its component models or accuracy of its confidence metric $A$.
@@
There are many approaches to approximating LLM confidence. A large subset of these methods involves deriving the confidence from the log-probabilities of the model outputs. In the case of LLMs, these log-probabilities, or “logits”, reflect a distribution over tokens and the confidence can be estimated using some variation of entropy or maximum probability at the individual token or sequence level. Some techniques that have shown recent promise involve examining consistency across multiple stochastic or permuted generations (Huang et al., 2025), but the additional computation required makes these approaches poorly suited to efficiency optimizations. Other researchers have asked generative models to explicitly report their confidence in their output with mixed results (Kadavath et al., 2022; Pawitan and Holmes, 2024), but benchmarks in some domains have found self reported confidence to be very poor, especially among smaller models (Omar et al., 2024).
@@
Different layers in a language model capture different types of information and fulfill different roles in inference (Tenney et al., 2019). One method of identifying and promoting fact-based and truthful outputs from LLMs is to consider the likelihood of different tokens at different layers within the model, emphasizing the effect of layers most likely to inject knowledge (Chuang et al., 2024). The internal layers of a pretrained model inform output rather than generate output directly and may hold useful information without the same biases as the output layer (Ji et al., 2022). Other, similar techniques applying an output head over internal layers have been used in early-layer exit strategies (Teerapittayanon et al., 2017), improve interpretability (Zhao et al., 2024), and to better understand how the internal layers and mechanisms function in LLMs (He et al., 2024; Tighidet et al., 2024).
@@
In the APRICOT method of Ulmer et al. (2024), a small auxiliary language model is used in support of a larger LLM to approximate the confidence of the larger model based on an embedding of the question and the model’s output. They demonstrate strong calibration scores and superior identification of mistakes compared to logit based sequence likelihood and self-assessed confidence. Further, ablations in this study indicate that confidence can reasonably be estimated without model outputs, a potential boon for cascading systems where confidence prior to execution is more valuable for efficiency.
3. Methodology
Figure 1. Bi-Directional deferral system diagram.
3.1. Cascade Definition
@@
As noted in Sec 1, existing cascade work generally defines the deferral choice only in terms of one model. Here we define our cascade in a form similar to that of Varshney and Baral (2022) but in terms of confidence based on two adjacent models in the cascade.
@@
3.2. Backwards Confidence Estimation
The post-invocation small model confidence method uses internal representations to understand how an output was influenced by the different layers in the model, and uses them to represent how reliable the output is. The general idea of using internal representations for better calibrated probabilities has been applied in Beigi et al. (2024) and Azaria and Mitchell (2023). Our formulation below is similar to that of Chuang et al. (2024), applied there to improve factuality, and here adapted for classification tasks. We apply it to develop better representations for confidence.
@@ 3.2. Backwards Confidence Estimation
where $\mathcal{J}\subseteq\{0,1,\ldots,N-1\}$ is a set of candidate internal layers. The confidence representation $Q_{M}(x)$ will then consist of a vector of these pseudo-probabilities $q_{j}(y|x)$ for some $\mathcal{J}$, of length $|\mathcal{Y}|\times(|\mathcal{J}|+1)$, to be used as the basis for features in the deferral model of Sec 3.4. Using multiple internal players provides an indicator of confidence for use in the deferral process by capturing the depth at which each potential response became more or less likely and the stability of each candidate output through the model.
[…]
@@
where $H(\cdot)$ is Shannon entropy. Existing cascade models are then generally special cases of this, with most just using $g_{\text{maxP}}(\cdot)$ (i.e., the maximum score from the final layer, and comparing to a threshold), noted by Gupta et al. (2024) as being widely used as it is “remarkably hard to surpass in most natural settings”.
3.3. Forwards Confidence Estimation
Ideally, we would be able to derive the same backwards confidence representation for the large model. However, as noted by Jitkrittum et al. (2023), in this kind of cascading scenario we do not have access to the large model to obtain these. Our idea, then, is to predict values that can take the place of these.
With the goal of calibrating LLMs in a QA context using their generations only, Ulmer et al. (2024) presented a method for setting confidence targets and training an additional model that predicts an LLM’s confidence based on its textual input and output alone. In our scenario, we do not have access to large model output, but do have access to the input; the ablation analysis in Appendix A.5 of Ulmer et al. (2024) suggests that input text alone could be feasible as the input to the additional predictor model.
Our pre-invocation large model confidence method thus uses a small auxiliary support model to learn the relationship between an input sequence and the expected large model output confidence. In order to train this model, we first produce a dataset by running once-off inference on a subset of the task data using the large model $M_{L}$. Following Zadrozny and Elkan (2001), entropy is calculated over the outputs of the model and samples are divided into bins by their entropy rank. Mean accuracy of samples in each bin provides a mapping from entropy to direct calibration target, then mean bin accuracy for each sample is used as the training target for the auxiliary forward confidence estimator.
@@ 3.3. Forwards Confidence Estimation
[…]
3.4. Deferral Model
@@
Since the representation of meaning throughout the model is itself complex and not yet fully understood (Jin et al., 2024), we feed all of the extracted pseudo-probabilities along with the forward confidence estimate to a deferral model $M_{D}$, which aims to identify samples that are most likely to be potential gains through deferral. Since the inputs to this model are numeric rather than text, traditional classifiers are adequate with minimal computational cost. We first define the ‘gain’ binary target, $y_{G,i}$, as samples where the output of small model $M_{S}$ does not match the true label and the output of large model $M_{L}$ does match the true label in the once-off generated training data:
$y_{G,i}=(M_{S}(x_{i})\neq y_{i})\land(M_{L}(x_{i})=y_{i})$
@@
Since the delegator model produces probabilistic outputs, we can trivially apply a threshold $10<\tau<$ 1 at which potential gains should be delegated to a larger model $M_{L}$ if $M_{D}>\tau$ (Algo 1).
4. Experimental Setup
4.1. Language Models
The publicly available pre-trained LLaMA-2 models are used as the cascading LLMs (Touvron et al., 2023), with the 13 billion parameter version used as the smallest cascade model, $M_{S}$, and the 70 billion parameter LLaMA-2 model as the large model, $M_{L}$. These models are open and accessible enough to conduct relevant experiments while being modern in architecture and competitive in performance. The primary models do not undergo any training or fine-tuning and are used as-is.
The forward confidence estimator is a pre-trained DeBERTa V3 model with a regression head at the base size with 86 million parameters (He et al., 2023). This model is approximately 0.7% the size of the smaller cascade model, resulting in a tiny increase in computation for the system.
The deferral model is a simple Random Forest Classifier using the Scikit Learn library (Pedregosa et al., 2011), trained on pseudo-probabilities from the backwards confidence estimator and the output of the forward confidence estimator to identify samples where the larger model will provide a better answer than the current output.
4.2. Datasets
The methods are applied to four commonly used multiple choice question-answering (MCQA) language datasets. We use BoolQ (Clark et al., 2019) yes/no question answering with context as our easiest task, and Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021), AI2 Reasoning Challenge (ARC) Science exam question dataset (Easy and Challenge subsets; ARC(E) and ARC(C) resp.) (Clark et al., 2018), and Common Sense Question Answering (CSQA) (Talmor et al., 2019) as more difficult benchmarks.
4.3. Cascade Models
Our primary model of interest is the one described in Sec 3, using the full features defined by $g_{\text{ident}}$ in the backwards confidence for the small model. We refer to this as the bi-directional model Bi-Dir.

### v2

@@ Abstract
[…]
123
1 Introduction
@@
In the last decade large, transformer-based, pretrained language models (LLMs) have surged in both popularity and capability, radically improving the state of the art in a wide range of tasks, from traditional Natural Language Processing (NLP) tasks to more general problems that were previously impossible, such as complex logical reasoning and the application of expert knowledge. However, these abilities have often required increasingly large model sizes where inference is expensive, raising concerns ranging from user budgets [6] to environmental impacts [40, 38], including among the NLP community [1]. In response to this, a diverse range of techniques have been developed to accelerate or improve the efficiency of LLMs.
@@
Some approaches aim to reduce the size of individual models like quantization [53] and model pruning [41]. Others have drawn from the field of adaptive computation, developing systems of models that collaborate to improve some aspect of performance. Conceptually relying on the fact that not all tasks are equally difficult, these approaches make the inference process dependent on the individual sample, leveraging smaller models or subsystems to handle easy instances. This concept has been used to provide large improvements in model throughput using draft-and-verify paradigms [5, 25] as well as to optimize the accuracy-efficiency trade-off [31]. Early-exit model cascading is a technique that has been used in other machine learning domains for some time [49, 34] that takes advantage of this principle by engaging a sequence of models of increasing size and capability. Inference is performed on each sample starting with the smallest model and exiting the cascade if the output of any model appears sufficient by some measure of confidence. This approach has recently been applied to NLP and LLMs and has delivered significant efficiency improvements [54]. However, there are some notable limitations that we address in this paper.
@@
In cascading models, there is typically a sequence of models of increasing complexity, $M_{1}\ldots M_{K}$, and Jitkrittum et al. [21] observe that standardly, the decision to defer from a smaller model $M_{i}$ to larger model $M_{i+1}$ is based only on some measure of confidence from the smaller model $M_{i}$. This can result in various failure modes, for example in confidence-based deferral erroneously forwarding samples where $M_{i+1}$ performs worse than $M_{i}$. Jitkrittum et al. [21] carry out a theoretical analysis for the two-model case that defines oracles that include probabilities from the larger model, formally characterising those cases where deferrals based on small-model probabilities alone are insufficient or non-optimal. For practical use, they define post-hoc estimators of their oracle rules that include theoretical confidence of both small and large models. Importantly, they note that since querying the larger model defeats the entire purpose of cascades, they use only the outputs of the small model in their estimators. Their experiments on computer vision datasets, however, do not show much improvement over the standard use of small model confidence.
@@
The first core idea of the present paper is to use a proxy for large model confidences that can be obtained without querying the large model, by building a much smaller model whose only role is to predict large model confidences. For this, we draw on recent work on calibrating LLMs by Ulmer et al. [47] that will allow us to use only model inputs. We then have ‘backward’, post-invocation confidence for the smaller model, and a ‘forward’, pre-invocation ’proxy confidence’ estimation for larger models using a small auxiliary model.
@@
There is then the question of how to relate these two kinds of confidence estimation. We cannot expect that they will be calibrated with respect to each other: in the first place, one is directly taken from the inference model and the other from a proxy; in addition, it is known that larger models produce better calibrated results than smaller models [57]. Our second core idea is then to use a meta-model for making deferral decisions in a cascade, extending their use from single-model calibration [7, 39]. This then requires a richer representation of features for model confidence: where existing cascading work largely uses final-layer maximum softmax probability output or entropy, we use intermediate layer probing results, which have been shown to be effective in improving calibration in various contexts [33, 29, 8].
@@
  • We present a new bi-directional approach to cascading with a meta-model incorporating (i) richer representations of multiple model confidences using internal model states, and (ii) novel proxy confidence measures that avoid queryring large models.
  • We find the bi-directional approach outperforms methods that use only smaller model confidence. We also find that, by itself, using internal state representations for smaller model confidence improves over the standard use of final layer maximum probability.
@@ 2 Related Work
The accuracy-efficiency trade-off is a pervasive concern in ML, particularly in the field of LLMs, and there are many methods for reducing the computational requirements of models while minimally compromising on output quality [23]. The idea of making an early-exit from inference is one that has existed for many years [49], but has seen renewed interest with the increasing size and commercial viability of modern models [37]. Broadly this involves managing the interaction between each sample and the inference system to minimize wasteful computation. This can been applied at the model-level, directing the flow of samples through a set of discrete models, or at the layer-level, controlling the number of dependent layers to be used within a single neural network model [28].
2.1 Model Cascades
@@
In early-exit model cascades inference is first performed on a sample using the smallest available model, then some estimate of the confidence of the result is used to determine whether the sample should be deferred to the next largest model or the output should be accepted as is. This process occurs iteratively until the output is accepted or the largest model in the cascade is invoked [48]. The decision to accept or defer is typically a comparison between the confidence and some static or dynamic threshold, which can be adjusted to satisfy the desired trade-off [51, 24].
@@
Approaches in existing literature vary in number and type of models in the cascade. Lebovitz et al. [24] find that two model cascades perform well, with three model cascades performing marginally better at the cost of increased system complexity, while others [6] include many models to take advantage of disparate commercial API prices. Nie et al. [31] cascade 3 heterogeneous models (Logistic Regression, BERT and GPT/LLaMA) with online optimization, treating the largest model outputs as ground-truth with strong results; however, this constrains the application to tasks where traditional models (LR) are viable and assumes the success of largest available model. Choice of confidence metric $A$ also varies, with maximum probability and entropy both being frequently used and effective with some conflicting results regarding which is better [24, 51]. Our work broadly fits into this part of the space, following the observations of Jitkrittum et al. [21] about the potential (but not necessarily realised) value of using confidence from multiple models in a cascade.
@@
Other recent work has focussed on LLM-specific features. To note three: Gupta et al. [14] use language model cascades to accomplish generative tasks; Yue et al. [54] use Chain-of-Thought prompting to improve small model confidence; [50] alter the LLM training paradigm to produce models best suited to cascading.
A task related to model cascading is Learning to Defer (L2D) [26]; Jitkrittum et al. [21] discusses their relationship. In summary, most work in the L2D space assumes a single stage framework involving training models from scratch as part of simultaneous learning of predictor and deferral function, which is not feasible for LLMs Mao et al. [27], Narasimhan et al. [30]. More generally, work in the L2D space is aimed at developing a framework in which models can be trained to make a prediction or defer to a (usually human) expert, using a combined cost function including some expert invocation penalty. Narasimhan et al. [30] note that L2D and model cascading can be similar under some constraints; however, traditional cascading approaches like ours, model only the expected performance improvement of deferral, allowing the user to determine the cost-performance tradeoff in a post-hoc manner without retraining.
2.2 LLM Calibration
@@
A model’s calibration is its ability to express some level of confidence associated with an output that accurately reflects the likelihood that it is correct [13]. LLM calibration generally has great value to researchers and industry applications, and is the subject of significant research efforts [12]. The acceptance of LLMs in real-world application depends on their trustworthiness and reliability [4], while in cascading context, calibration has direct and measurable benefits on performance. The success of any model cascade depends on the quality of its deferral process, and therefore the calibration of its component models or accuracy of its confidence metric $A$.
@@
There are many approaches to approximating LLM confidence. A large subset of these methods involves deriving the confidence from the log-probabilities of the model outputs. In the case of LLMs, these log-probabilities, or “logits”, reflect a distribution over tokens and the confidence can be estimated using some variation of entropy or maximum probability at the individual token or sequence level. Some techniques that have shown recent promise involve examining consistency across multiple stochastic or permuted generations [18], but the additional computation required makes these approaches poorly suited to efficiency optimizations. Other researchers have asked generative models to explicitly report their confidence in their output with mixed results [22, 35], but benchmarks in some domains have found self reported confidence to be very poor, especially among smaller models [32].
@@
Different layers in a language model capture different types of information and fulfill different roles in inference [44]. One method of identifying and promoting fact-based and truthful outputs from LLMs is to consider the likelihood of different tokens at different layers within the model, emphasizing the effect of layers most likely to inject knowledge [8]. The internal layers of a pretrained model inform output rather than generate output directly and may hold useful information without the same biases as the output layer [19]. Other, similar techniques applying an output head over internal layers have been used in early-layer exit strategies [43], improve interpretability [56], and to better understand how the internal layers and mechanisms function in LLMs [15, 45].
@@
In the APRICOT method of Ulmer et al. [47], a small auxiliary language model is used in support of a larger LLM to approximate the confidence of the larger model based on an embedding of the question and the model’s output. They demonstrate strong calibration scores and superior identification of mistakes compared to logit based sequence likelihood and self-assessed confidence. Further, ablations in this study indicate that confidence can reasonably be estimated without model outputs, a potential boon for cascading systems where confidence prior to execution is more valuable for efficiency.
3 Methodology
Figure 1: Bi-Directional deferral system diagram.
3.1 Cascade Definition
@@
As noted in Sec 1, existing cascade work generally defines the deferral choice only in terms of one model. Here we define our cascade in a form similar to that of Varshney and Baral [48] but in terms of confidence based on two adjacent models in the cascade.
@@
3.2 Backwards Confidence Estimation
The post-invocation small model confidence method uses internal representations to understand how an output was influenced by the different layers in the model, and uses them to represent how reliable the output is. The general idea of using internal representations for better calibrated probabilities has been applied in Beigi et al. [3] and Azaria and Mitchell [2]. Our formulation below is similar to that of Chuang et al. [8], applied there to improve factuality, and here adapted for classification tasks. We apply it to develop better representations for confidence.
@@ 3.2 Backwards Confidence Estimation
where $\mathcal{J}\subseteq\{0,1,\ldots,N-1\}$ is a set of candidate internal layers. The confidence representation $Q_{M}(x)$ will then consist of a vector of these pseudo-probabilities $q_{j}(y|x)$ for some $\mathcal{J}$, of length $|\mathcal{Y}|\times(|\mathcal{J}|+1)$, to be used as the basis for features in the deferral model of Sec 3.4. Using multiple internal layers provides an indicator of confidence for use in the deferral process by capturing the depth at which each potential response became more or less likely and the stability of each candidate output through the model.
We then apply some function $g$ to $Q_{M}$ to derive the backward features for the small model $\text{Conf${}_{B}$}(M_{S}(x))=g(Q_{M_{S}}(x))$ (Algo 1) used in our deferral model (Sec 3.4). The specific functions we consider in this paper are:
@@
where $H(\cdot)$ is Shannon entropy. Existing cascade models are then generally special cases of this, with most just using $g_{\text{maxP}}(\cdot)$ (i.e., the maximum score from the final layer, and comparing to a threshold), noted by Gupta et al. [14] as being widely used as it is “remarkably hard to surpass in most natural settings”.
3.3 Forwards Confidence Estimation
Ideally, we would be able to derive the same backwards confidence representation for the large model. However, as noted by Jitkrittum et al. [21], in this kind of cascading scenario we do not have access to the large model during inference to obtain these. Our idea, then, is to predict values that can take the place of these.
With the goal of calibrating LLMs in a QA context using their generations only, Ulmer et al. [47] presented a method for setting confidence targets and training an additional model that predicts an LLM’s confidence based on its textual input and output alone. In our scenario, we do not have access to large model output, but do have access to the input; the ablation analysis in Appendix A.5 of Ulmer et al. [47] suggests that input text alone could be feasible as the input to the additional predictor model.
Our pre-invocation large model confidence method thus uses a small auxiliary support model to learn the relationship between an input sequence and the expected large model output confidence. In order to train this model, we first produce a dataset by running once-off inference on a subset of the task data using the large model $M_{L}$. Following Zadrozny and Elkan [55], entropy is calculated over the outputs of the model and samples are divided into bins by their entropy rank. Mean accuracy of samples in each bin provides a mapping from entropy to direct calibration target, then mean bin accuracy for each sample is used as the training target for the auxiliary forward confidence estimator.
@@ 3.3 Forwards Confidence Estimation
The forward feature for the large model is then $\text{Conf${}_{F}$}(M_{L}(x))=M_{A}(x)$.
3.4 Deferral Model
@@
Since the representation of meaning throughout the model is itself complex and not yet fully understood [20], we feed all of the extracted pseudo-probabilities along with the forward confidence estimate to a deferral model $M_{D}$, which aims to identify samples that are most likely to be potential gains through deferral. Since the inputs to this model are numeric rather than text, traditional classifiers are adequate with minimal computational cost. We first define the ‘gain’ binary target, $y_{G,i}$, as samples where the output of small model $M_{S}$ does not match the true label and the output of large model $M_{L}$ does match the true label in the once-off generated training data:
$y_{G,i}=\begin{cases}1&\text{if }(\mathsf{MS}(x_{i})\neq y_{i})\land(\mathsf{ML}(x_{i})=y_{i})\\ 0&\text{otherwise}\end{cases}$
@@
Since the deferral model produces probabilistic outputs, we can trivially apply a threshold $0<\tau<1$ at which potential gains should be deferred to a larger model $M_{L}$ if $M_{D}>\tau$ (Algo 1).
4 Experimental Setup
4.1 Models
The publicly available pre-trained LLaMA-2 models are used as the cascade component LLMs [46], with the 13 billion parameter version used as the smallest cascade model, $M_{S}$, and the 70 billion parameter LLaMA-2 model as the large model, $M_{L}$. These models are open and accessible enough to conduct relevant experiments while being modern in architecture and competitive in performance. The primary models do not undergo any training or fine-tuning and are used as-is.
The forward confidence estimator is a pre-trained DeBERTa V3 model with a regression head at the base size with 86 million parameters [16]. This model is approximately 0.7% the size of the smaller cascade model, resulting in a tiny increase in computation for the system.
The deferral model is a simple Random Forest Classifier using the Scikit Learn library [36], trained on pseudo-probabilities from the backwards confidence estimator and the output of the forward confidence estimator to identify samples where the larger model will provide a better answer than the current output.
All model training uses a 5-fold cross-validation approach and only unseen test fold results are reported.
4.2 Datasets
The methods are applied to four commonly used multiple choice question-answering (MCQA) language datasets. We use BoolQ [9] yes/no question answering with context as our easiest task, and Massive Multitask Language Understanding (MMLU) [17], AI2 Reasoning Challenge (ARC) Science exam question dataset (Easy and Challenge subsets; ARC(E) and ARC(C) resp.) [10], and Common Sense Question Answering (CSQA) [42] as more difficult benchmarks.
4.3 Cascade
Our primary method of cascade deferral is using the outputs of the model described in Sec 3.4, using the full features defined by $g_{\text{ident}}$ in the backwards confidence for the small model combined with $\text{Conf}_{F}$. We refer to this as the bi-directional model Bi-Dir.